OpenStudy (anonymous):

Find the slant of a cone with surface area 103.67 square feet and radius 3 feet.

hartnn (hartnn):

surface area of cone is \(\Large \pi r (r+l)\) r is the radius, l is the slant height use this formula

OpenStudy (anonymous):

l=(r^2+h^2)^1/2

OpenStudy (anonymous):

u can calculate h with the help of surface area...

OpenStudy (anonymous):

and l is ur slant height....

hartnn (hartnn):

\(103.67 \approx 33\pi \) r=3 \(33\pi = 3\pi (3+l)\) only l is unknown :)

OpenStudy (anonymous):

Thanks so much! I wasn't sure what formula to use. My final answer was h = 8 feet.

hartnn (hartnn):

\(\huge \checkmark \)